The area of a square is 81 square centimeters. Find the length of a side. Find the length of the diagonal.
Answer by mananth(16949)

area of square = 81
side = sqrt(81)
side = 9
diagonal^2 = (9)^2+(9)^2
diagonal = sqrt(162)
diagonal = 9sqrt(2)